Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 2.4 h, and Car B traveled the distance in 4 h. Car A traveled 22 mph faster than Car B.

How fast did Car A travel?

Answer by richwmiller(17219)

2.4*r=4*(r-22)
2.4*r=4*r-88
88=4*r-2.4*r
88-1.6r
88/1.6=r
880/16=r
110/2=r
r=55
r-22=33